Twenty-five feet of copper tubing weights 14 pounds. How much will 140 feet of the same tubing weigh?
Answer by checkley75(3666)

25/14=140/x cross multiply
25x=14*140
25x=1960
x=1960/25
x=78.4 pounds for the 140 feet of copper tubing.